# FedAGHN: Personalized Federated Learning with Attentive Graph HyperNetworks

### v1

@@ C.1 Data Heterogeneity
In both settings, we divide the local dataset into a training set, a validation set, and a test set for each client after distributing the whole dataset. The data split ratio for training, validation, and test is set to 7:1:2.
C.2 Baselines and Evaluation
@@
C.3 Implementation Details
@@
For the 4-layer CNN architecture with BN layers, we train it for 200 global communication rounds (1000 iterations). For ResNet-18, we train it for 80 rounds (400 iterations) on CIFAR100 and 40 rounds (200 iterations) on Tiny-Imagenet to ensure the empirical convergence.
[…]
All experiments are repeated for three times, and we report the mean and standard deviation.
Appendix D Supplementary of Experimental Results
D.1 Ablation Study
To verify the effectiveness of designed components in FedAGHN, ablation experiments are performed in the practical setting. We consider trainable parameters $p$ and $q$ in AGHN, together with their trainability (T) and layer-wise differences (L). The results of the ablation experiments are shown in Table 1, where ‘with’ is abbreviated as ‘w/’ and ‘without’ is abbreviated as ‘w/o’.
FedAGHN performs best when all its designed components are fully considered, reflecting the positive contribution of FedAGHN’s designed components to its performance. Specifically, first of all, the most severe performance decrease occurs when $p$ and $q$ are absent, highlighting the significance of FedAGHN in explicitly tuning collaborative relationships. Secondly, freezing $p$ and $q$ during the FL process leads to performance decrease, which demonstrates that AGHN’s adaptive optimization of trainable parameters can better tune current collaborative relationships. Thirdly, removing $p$ and $q$ separately results in a greater decrease in performance without $p$ than without $q$, indicating the importance of tuning the client’s self-weighting value and adaptive optimization. Finally, performance also decreases when there is no layer-wise difference in $p$ and $q$, highlighting the importance of layer-wise collaboration tuning.
[…]
@@ D.1 Ablation Study
D.2 Different Degrees of Heterogeneity
@@
A smaller $\beta$ indicates higher data heterogeneity as illustrated in Figure 1. The experimental results of FedAGHN and baseline methods under various degrees of data heterogeneity are presented in Table 2 of the main manuscript, indicating that some personalized aggregation-based methods show lower performance, such as CFL, FedAMP and H-FedAMP, while fine-grained personalized aggregation-based methods such as pFedLA and FedALA perform well. Furthermore, pFedLA and FedPAC show competitive performance, possibly because both of them partially address the personalization needs of different model layers.
@@
Appendix E Supplementary Analysis of Collaboration in FedAGHN
E.1 Case Study Setting
We randomly sample $s$of the data (80% by default) and divide it into 5 groups based on the classes of labels. The $N=20$ clients are divided into 5 groups, and each group is assigned data from the same classes. The remaining 1-$s$of the data is distributed to all clients in an IID manner.
E.2 Analysis of Collaboration Graphs
The averaged collaboration weights are significantly different for clients with different relationships in FedAGHN, where the self-weighting value is smaller and the weights for clients within the same group are larger. It might be because clients in the same group have more diverse information. Furthermore, CFL maintains the same weight for clients within a group, and H-FedAMP assigns slightly different weights for clients with different relationships.
Additionally, we find that setting the self-weighting value to a relatively smaller value leads to better performance in our experiments. This may be due to the fact that FedAGHN provides detailed descriptions of collaborative relationships. Assuming effective capture of collaborative relationships, reducing the self-weighting value facilitates the acquisition of more accurate and diverse information.
E.3 Analysis of Dynamic Collaboration
To analyze the dynamic collaboration in FedAGHN, Figure 2 illustrates the evolution of collaboration weights among different methods.
@@ E.3 Analysis of Dynamic Collaboration
The evolution of collaboration weights for CFL is dependent on clustering, where similar clients are given high weights but the collaboration of other clients is completely ignored. The weight differences among collaborating clients in H-FedAMP are slight, and the self-weighting value remains constant throughout the training process.
To analyze the evolution trends of the collaborative relationships for different model layers, Table 4 of the main manuscript shows the changes of weights for the whole model, shallow layers and deep layers of model in FedAGHN. The definition of different stages follows the description in Figure 2 and the definition of ‘Self’, ‘Similar’ and ‘Other’ follows the explanation in Section E.2. It can be observed that FedAGHN fully considers the differences in collaborative requirements across different model layers.

### v2

@@ 4.2 Data Heterogeneity
For both settings, we divide the local dataset into a training set, a validation set, and a test set for each client after distributing the whole dataset. The data split ratio for training, validation, and test is set to 7:1:2.
4.3 Baselines and Evaluation
@@
For each method, we save the best model that achieves the best performance on the local validation set during the training process, and evaluate it on the local test set of each client. The average of test accuracy across all clients is calculated. All experiments are repeated for three times, and the results are reported in mean $\pm$ standard deviation.
Settings Pathological setting Practical setting (Dir(0.1))
CIFAR10 CIFAR100 $\text{CIFAR100}^{*}$ TINY CIFAR10 CIFAR100 $\text{CIFAR100}^{*}$ TINY
4.4 Implementation Details
@@
In the specific experimental implementation, we use the SGD optimizer and set the local batch size to 64. The number of local training epochs is set to 5, and the local learning rate for models is set to 0.01. We consider 20 clients participating with the join rate of 1. For the 4-layer CNN architecture with BN layers, we train it for 200 global communication rounds (1000 iterations). For ResNet-18, we train it for 80 rounds (400 iterations) on CIFAR100 and 40 rounds (200 iterations) on Tiny-Imagenet to ensure the empirical convergence.
For the proposed FedAGHN, the learning rate of the hypernetwork is set to $\eta_{hn}=0.005$. We tune $p_{i}$ from $\{0.02,0.03,0.06\}$ and $q_{i}$ from $\{0.1,0.2,0.5,1,3,10\}$ for different datasets and Non-IID settings. For the other methods, we generally follow Ye et al. (2023) to tune and set the hyper-parameters for these baselines [^0] For FedProx and FedProx-FT, we set the regularization weighted hyperparameter $\lambda=0.01$. For Ditto, we use $\lambda=1$ and $s=5$. For FedRep, we set $\tau=5$. For pFedHN, we set $\eta_{hn}=0.005$. For FedCP, we set $\lambda=5$ on 4-layer CNN architecture with BN layers and $\lambda=1$ on ResNet-18. For FedPAC, we set $\lambda=1$. For CFL, we set $\epsilon_{1}=0.5$, $\epsilon_{2}=0.5$ on 4-layer CNN architecture with BN layers and $\epsilon_{1}=0.4$, $\epsilon_{2}=1.6$ on ResNet-18, set $\text{minimum cluster size}=4$ on case study setting and $\text{minimum cluster size}=2$ on other settings, set $\text{start clustering round}=20/10/8$ for $\text{global rounds}=200/80/40$. For FedAMP, we set $\alpha_{K}=100$, $\lambda=1$, $\sigma=10000$. And for H-FedAMP, we set $\sigma=100$, $\omega_{ii}=0.05$, $\lambda=1$, $\alpha_{K}=1$ on 4-layer CNN architecture with BN layers and $\omega_{ii}=0.1$ on ResNet-18. For pFedLA, we set $\eta_{hn}=0.005$, $\text{embedding dimension}=100$ and $\text{hidden layer dimension}=100$. For FedALA, we follow the setting as original paper..
5 Results and Discussions
  • RQ2: Can and how the designs in FedAGHN contribute positively to performance? (Section 5.2 and Section 5.5) (Table 4 and Figure 5)
  • RQ3: Is FedAGHN still effective across various degrees of statistical heterogeneity as well as other scenarios? (Section 5.3 and Section 5.4) (Table 5 and Table 6)
  • RQ4: Does the lightweight design of FedAGHN reduce the additional system overhead introduced by employing hypernetworks? (Section 5.6) (Table 7)
5.1 Main Experiment
To verify the superiority of the proposed FedAGHN against the 15 baseline methods, we fairly test them over the aforementioned datasets and two different Non-IID settings. The best and second best performance are bolded and underlined, respectively, in Table 3.
Overall, FedAGHN outperforms almost all baselines and achieves the state-of-the-art performance for all cases across the two Non-IID scenarios. This might have been due to FedAGHN benefiting from its full consideration of personalization at the layer-wise level and its adaptively tuned client collaborative relationships. The only case where FedAGHN does not achieve the first place is in the pathological setting of CIFAR10, where it takes the second place, narrowly trailing behind FedPAC. This may have been due to FedPAC making it easier to achieve feature alignment when both the number of model parameters and the variety of data sample labels are small.
The traditional FL methods FedAvg and FedProx perform worse than local training, whereas the local adaptation methods FedAvg-FT and FedProx-FT, remain competitive among various PFL approaches.
With regaed to the personalized aggregation-based methods, CFL performs competitively in the practical setting of TINY but performs worse in the other settings. This might have been due to the tendency of CFL to collaborate with a few clients whose updates of personalized local model parameters are most similar, while ignoring the general information from the other clients, making it less adaptable to different scenarios. On the other hand, H-FedAMP achieves high performance in the pathological setting of TINY and it outperforms FedAMP in most scenarios, possibly owing to its calculation of relative collaboration weights for each collaborating client, and its more flexible calculation of collaboration weights than that offered by FedAMP.
This study finds that finer-grained aggregation may result in improved performance. pFedLA performs well in most cases, possibly owing to its fine-grained personalized aggregation. However, pFedLA is not adaptable to different scenarios, possibly because of the reliance of its weight generation relies entirely on the optimization effects of the hypernetworks model parameters and the clients embeddings. FedALA performs worse than pFedLA in most cases, because it integrates information from both local and global models at a finer granularity but overlooks the selection of collaborative relationships with other clients. By comparison, our proposed FedAGHN outperforms all the personalized aggregation-based methods tested herein, possibly as a result of its ability to explicitly model collaborative relationships and adaptively tune collaboration weights while performing finer-grained aggregation.
Design choices in FedAGHN $p$ $q$ T L CIFAR10 CIFAR100 $\text{CIFAR100}^{*}$ TINY
@@ 5.1 Main Experiment
5.2 Ablation Study
To verify the effectiveness of the designed components in FedAGHN, ablation experiments are performed in the practical setting. We consider trainable parameters $p$ and $q$ in AGHN, together with their trainability (T) and layer-wise differences (L). The results of the ablation experiments are shown in Table 4, where ‘with’ is abbreviated as ‘w/’ and ‘without’ is abbreviated as ‘w/o’.
FedAGHN performs best when all of its designed components are fully considered, reflecting the positive contribution of each of the designed components to the overall performance. Specifically, first of all, the most severe performance decrease occurs when $p$ and $q$ are absent, highlighting the significance of FedAGHN in explicitly tuning collaborative relationships. Second, freezing $p$ and $q$ during the FL process leads to performance decrease, which demonstrates that AGHN’s adaptive optimization of trainable parameters can better tune current collaborative relationships. Third, removing $p$ and $q$ separately results in a greater decrease in performance without $p$ than without $q$, indicating the importance of tuning the client’s self-weighting value and adaptive optimization. Finally, performance also decreases when there is no layer-wise difference in $p$ and $q$, highlighting the importance of layer-wise collaboration tuning.
5.3 Different Degrees of Heterogeneity
@@
To investigate the performance of FedAGHN under various degrees of data heterogeneity, we manipulate the parameter $\beta$ in Dir($\beta$) to generate heterogeneous data with different distributions on the CIFAR10 dataset. A smaller $\beta$ indicates higher data heterogeneity as illustrated in Figure 4. The experimental results of FedAGHN and the baseline methods under various degrees of data heterogeneity are presented in Table 5.
Settings Dir(0.1) Dir(0.2) Dir(0.5)
Local 85.00±0.42 80.47±0.21 71.41±0.12
FedAvg 60.72±1.13 60.04±1.52 65.06±0.11
FedAvg-FT 87.54±0.24 84.95±0.34 79.51±0.16
FedProx 60.58±0.84 60.02±1.28 65.05±0.08
FedProx-FT 87.65±0.34 84.69±0.33 79.60±0.33
Ditto 86.77±0.27 84.19±0.32 79.11±0.43
FedRep 87.27±0.09 84.22±0.31 77.47±0.34
pFedHN 83.60±0.42 80.83±0.34 73.18±0.41
FedCP 87.78±0.27 85.25±0.42 79.57±0.16
FedPAC 87.59±0.17 85.39±0.10 79.35±0.66
CFL 87.17±0.16 83.96±0.18 77.13±0.25
FedAMP 85.07±0.43 80.57±0.33 71.77±0.23
H-FedAMP 86.01±0.32 82.64±0.64 75.77±0.55
pFedLA 88.10±0.30 84.97±0.16 79.37±0.56
FedALA 87.63±0.18 84.89±0.21 79.64±0.36
FedAGHN 88.32±0.29 85.45±0.29 79.85±0.52
Table 5: The accuracy (%) on CIFAR10 for different heterogeneity.
@@
FedAGHN outperforms all baselines across various degrees of data heterogeneity, which might have been owed to the consideration of fine-grained aggregation and personalization at layer-wise. It is evident that some of the personalized aggregation-based methods, such as CFL, FedAMP and H-FedAMP, yield show lower performance, whereas the fine-grained personalized aggregation-based methods such as pFedLA and FedALA perform well. Furthermore, pFedLA and FedPAC show competitive performance, possibly as a result of their capability to partially address the personalization needs of different model layers.
5.4 Performances on Other Scenarios
In addition, we have conducted the following experiments to account for a broader range of realistic scenarios.
Extreme heterogeneity: To simulates a scenario of extreme heterogeneity in label distribution, we divide the CIFAR-100 dataset into 20 independent clients, each client is randomly assigned local data from 5 distinct labels, with no label overlap across clients (i.e., each client has an exclusive set of 5 labels, ensuring completely disjoint label distributions).
Cross-device FL: To verify the performance under the cross-device FL scenario with more client participation, we expand the number of clients to 50/100 in the pathological heterogeneous scenario of the CIFAR-100 dataset, while maintaining the other experimental settings consistent with the main experiment. To ensure that the amount of data in each category for each client is consistent, we assign 20/8/4 classes to each client when the number of clients is 20/50/100, respectively.
The results for these two additional scenarios are summarized as Table 6. The experimental results show that our method can still maintain the superior performance under this extremely heterogeneous scenario and with an even larger number of client participation, consistent with the existing experimental conclusions.
Datasets Extreme Num=20 Num=50 Num=100
Local 81.53±0.44 52.88±0.45 66.98±0.53 78.00±0.10
CFL 83.30±0.07 58.36±0.35 70.38±0.34 80.21±0.06
H-FedAMP 80.43±0.10 55.47±0.49 66.79±0.49 77.41±0.30
FedAGHN 83.46±0.22 61.27±0.14 73.63±0.17 80.69±0.33
Table 6: The accuracy (%) of compared methods in case study
5.5 Sensitivity Analysis on p and q
For the proposed FedAGHN, for each client i, we tune $p_{i}$ from $\{0.02,0.03,0.06\}$ and $q_{i}$ from $\{0.1,0.2,0.5,1,3,10\}$ across different datasets and Non-IID settings. The initial values of $p_{i}$ and $q_{i}$ are applied uniformly to all R layers of the model.
Figure 5: Sensitivity analysis on p and q using the CIFAR-10 dataset under the Dir(0.1) partition setting.
To further explore the roles of $p_{i}$ and $q_{i}$ in our method, we conduct a sensitivity analysis on $p_{i}$ and $q_{i}$ using CIFAR-10 under the Dir(0.1) partition setting. The results are presented in Figure 5. The selections of $p_{i}$ and $q_{i}$ are as following:
  1. 1.
    $p_{i}$ represents the self-participation strength of client i. When the number of clients is 20, the average weight is approximately 0.05. Therefore, we set the initial range of $p_{i}$ to $\{0.02,0.03,0.06\}$ to reflect different degrees of self-participation.
  2. 2.
    $q_{i}$ serves as a scaling coefficient that adjusts the variability of collaboration weights (which can also be interpreted as the reciprocal of a temperature coefficient). Accordingly, $q_{i}$ is tuned from $\{0.1,0.2,0.5,1,3,10\}$ to capture varying levels of weight differentiation (or smoothness).
Overall, FedAGHN exhibits low sensitivity to these parameters. While different combinations of $p_{i}$ and $q_{i}$ introduce some variation in performance, the accuracy difference remains within approximately 0.8%. This robustness can be attributed to the dynamic optimization strategy and the integration of prior knowledge within AGHN’s tunable attention mechanism. Although the selection of $p_{i}$ and $q_{i}$ in the tunable attention mechanism introduces some variability, the dynamic optimization and prior knowledge substantially enhance the overall robustness.
5.6 System Overhead
To assess the effectiveness of FedAGHN’s lightweight design for hypernetworks, Table 7 presents a comparative analysis of the system overhead incurred by PFL methods employing hypernetworks.
Communication: Each client incurs a communication cost of $S$ for transmitting its local model. All PFL methods employing hypernetworks involve no additional communication overhead, as the hypernetworks are maintained on the server side.
Storage: For server-side storage cost (maximum required space during computation), $M$ denotes the local model parameter size, with $N$ clients participating. $M_{\text{hn}}$ represents storage overhead for maintaining hypernetwork. $M_{\text{hn}}$ varies across algorithms (pFedHN, pFedLA and FedAGHN for 1.13B, 115K and 82 trainable parameters). FedAGHN, which focuses on the design of critical parameters, has the smallest hypernetwork size. Thus, the model size $M$ is 11.23M, and the storage cost of $M_{\text{hn}}$ in FedAGHN is much smaller than that of $M$.
Computation: In terms of overall computation overhead and time during the aggregation phase, FedAGHN shows relatively low cost compared to other personalized aggregation methods, despite maintaining $N$ AGHNs. This efficiency benefits from the introduction of a small number of key parameters in the AGHN’s lightweight attention mechanism. $C$ denotes the computational cost of the weighted aggregation operation, and $C_{\text{hn}}$ represent the computational overhead of performing personalized aggregation using the hypernetwork. Set $N=20$ to evaluate time costs. The computation cost $C_{\text{hn}}$ of AGHN comprises three components: attention computation, on-graph aggregation, and hypernetwork update. The computation times for pFedHN, pFedLA, and FedAGHN are 1.77s, 1.54s, and 1.02s, respectively.
Scalability: For the scaling overhead, although FedAGHN maintains $N\times R$ collaboration graphs, its primary purpose is to enable personalized aggregation. Once the personalized aggregation is completed, the corresponding collaboration graph no longer needs to reside in memory (regardless of whether it is at the layer-wise or the client-wise). For example, after the $r$-th layer of client $i$ has completed personalized aggregation, $G_{i}^{r}$ can be safely discarded. Consequently, even when the number of clients $N$ scales to a very large value, this design prevents excessive memory overhead.
Algorithm Comm. Server Storage Computation Cost
pFedHN $S$ $M+M_{\text{hn}}$ $C_{\text{hn}}$
pFedLA $S$ $N\times M+N\times M_{\text{hn}}$ $N\times C_{\text{hn}}$
FedAGHN $S$ $N\times M+N\times M_{\text{hn}}$ $N\times C_{\text{hn}}$
Table 7: Comparison of system overhead for PFL methods employing hypernetworks
6 Analysis of Collaboration in FedAGHN
  • RQ5: Can the collaboration graphs in FedAGHN effectively capture the collaborative relationships among clients? (Section 6.2.1) (Figure 6, Table 8)
  • RQ6: Can FedAGHN meet the personalized requirements for layers and show differences among different layers? (Section 6.2.2) (Figure 7)
  • RQ7: Can FedAGHN dynamically tune collaborative relationships during FL process? (Section 6.3) (Figure 8)
  • RQ8: How do the collaborative relationships of different model layers captured by FedAGHN change throughout the FL process? (Section 6.3) (Table 10)
6.1 Case Study Setting
In this section, we establish a case study setting to investigate the collaborative relationship learned by FedAGHN. The Non-IID scenario is introduced following Xu et al. (2023). We randomly sample ss% of the data (80% by default) and divide them into 5 groups based on the classes of labels. The $N=20$ clients are divided into 5 groups, and each group is assigned data from the same classes. The remaining 1-ss% of the data is distributed to all clients in an IID manner.
Therefore, clients within each group follow the same data distribution, while clients across groups have different dominant classes. In this context, there are evident differences between clients within the same group and clients in different groups, and thus, we can intuitively estimate whether the collaborative relationships among clients are appropriately captured.
Table 8 has shown the performance of these methods in case study setting. FedAGHN performs best on these datasets in this setting, which could be explained by the appropriate collaboration graphs analyzed in Section 6.2 and 6.3. CFL has relatively high performance in most cases, possibly because CFL can collaborate with the clients in the same group, but it does not collaborate with the clients in different groups at all, ignoring the information from clients with different data distributions. H-FedAMP assigns smaller weights to the clients in different groups, but its performance is even worse than local training on $\text{CIFAR100}^{*}$, probably because the differences in weights among clients may not be sufficiently large.
Datasets CIFAR10 CIFAR100 $\text{CIFAR100}^{*}$ TINY
Local 73.96±0.10 40.25±0.25 28.56±0.40 14.99±0.04
CFL 79.66±0.21 49.55±0.34 38.59±0.46 25.93±0.29
H-FedAMP 78.23±0.49 46.11±0.67 27.09±0.81 26.74±0.22
FedAGHN 81.09±0.30 51.14±0.27 39.12±0.48 26.83±0.14
Table 8: The accuracy (%) of compared methods in case study
6.2 Analysis of Collaboration Graphs
6.2.1 Comparison with other methods
To evaluate the efficacy of personalized aggregation methods in capturing collaborative relationships, the visualization of collaboration graphs in CFL, H-FedAMP and FedAGHN are shown in Figure 6-a, 6-b and 6-c respectively.
As illustrated in Figure 6, FedAGHN can effectively capture collaborative relationships among clients, with small self-weighting value. Overall, the collaboration weights of FedAGHN generally align with the similarity of clients’ data distributions in the case study setting. Moreover, FedAGHN has clear weight differences both intra-group and inter-group. For the clients within a same group, there are significant weight differences among clients, which implies that higher weights are assigned to more beneficial clients. Besides, it is shown in Figure 6-a that all clients within a group are assigned the same collaboration weights in CFL, while clients outside the group do not participate in collaboration. Furthermore, H-FedAMP introduces a fixed self-weighting value, while assigning slightly higher weights to clients within the same group and lower weights to others.
@@ 6.2.1 Comparison with other methods
In FedAGHN, the averaged collaboration weights are significantly different for clients with different relationships, where the self-weighting value is smaller, whereas the weights for clients within the same group are larger. This might be because of the greater diversity of information among clients in the same group. On the other hand, CFL maintains the same weight for clients within a group, whereas H-FedAMP assigns slightly different weights for clients with different relationships.
6.2.2 Layer-wise collaboration graphs
To further investigate layer-wise personalization in FedAGHN, we visualize the collaboration graphs corresponding to different model layers on CIFAR100. Figure 7 reveals significant differences in the collaborative relationships captured by the shallow and deep layers of FedAGHN.
In the shallow layers (Figure 7-a, 7-b), the model tends to select more general information, with no significant differences between groups. FedAGHN introduces the tunable attentive mechanism to capture the fine-grained collaborative relationships among clients, which enables it to focus more on valuable information in the shallow layers, without excluding clients with different data distributions.
In the deep layers (Figure 7-c, 7-d), the model tends to select more specific information and prefers collaboration among clients within the same group. Significant differences can be observed between the groups, especially in the last model layer.
Additionally, we find that setting the self-weighting value to a relatively smaller value leads to better performance in our experiments. The reason behind this could be that FedAGHN provides the detailed descriptions of collaborative relationships. Under the assumption that the collaborative relationships are captured effectively, reducing the self-weighting value facilitates the acquisition of more accurate and diverse information.
6.3 Analysis of Dynamic Collaboration
To demonstrate that the proposed FedAGHN can dynamically tune the client collaboration during the training process, we record the changes of collaboration weights for different methods during the first, prior, middle, and final round of training. Figure 8 illustrates the evolution of collaboration weights among different methods.
@@ 6.3 Analysis of Dynamic Collaboration
Overall, the weight differences among collaborating clients continue to change during the training process, possibly indicating that the adaptive tuning of collaborative relationships may enhance FedAGHN’s performance. FedAGHN has a high demand for collaboration among clients with similar data distributions in the first round, which stabilizes in the middle round, but significant differences still exist among clients with different collaboration relationships. The self-weighting value gradually decreases during the FL process, indicating that clients prefer to collaborate with others. Therefore, FedAGHN’s adaptive tuning of collaborative relationships can better fulfill the collaboration requirements at different stages.
For comparison, the evolution of collaboration weights for CFL is dependent on clustering, where similar clients are given high weights but the collaboration with other clients is completely ignored. Meanwhile, the weight differences among the collaborating clients in H-FedAMP are slight, and the self-weighting value remains constant throughout the training process.
To analyze the evolution trends of the collaborative relationships for different model layers, Table 10 shows the changes of weights for the whole model, shallow layers and deep layers of model in FedAGHN. The definition of different stages follows the description in Figure 8 and the definition of ‘Self’, ‘Similar’ and ‘Other’ follows the explanation in Section 6.2.1. It can be observed that FedAGHN fully considers the differences in collaborative requirements across different model layers.
Client First Prior Middle Final
$\text{FedAGHN}_{\text{all}}$ Self 0.0291 0.0192 0.0136 0.0126
$\text{FedAGHN}_{\text{all}}$ Similar 0.1436 0.0758 0.0674 0.0733
$\text{FedAGHN}_{\text{all}}$ Other 0.0338 0.0471 0.0490 0.0480
$\text{FedAGHN}_{\text{shallow}}$ Self 0.0291 0.0228 0.0175 0.0155
$\text{FedAGHN}_{\text{shallow}}$ Similar 0.1270 0.0703 0.0598 0.0586
$\text{FedAGHN}_{\text{shallow}}$ Other 0.0369 0.0479 0.0502 0.0505
$\text{FedAGHN}_{\text{deep}}$ Self 0.0291 0.0155 0.0097 0.0097
$\text{FedAGHN}_{\text{deep}}$ Similar 0.1601 0.0814 0.0750 0.0880
$\text{FedAGHN}_{\text{deep}}$ Other 0.0307 0.0463 0.0478 0.0454
@@
7 Conclusion
This work proposed FedAGHN, a personalized aggregation-based method, to continuously capture appropriate collaborative relationships for fine-grained personalized aggregation. In each communication round, FedAGHN maintains a set of collaboration graphs for each client and each layer, which explicitly model the client-specific and layer-wise collaboration weights. The collaboration weights are obtained by the newly proposed tunable attentive mechanism. Finally, the personalized initial models are obtained via aggregating parameters over the learned collaboration graphs. Extensive experiments demonstrated the superiority of FedAGHN, which is attributed to its three important innovations (Layer-wise collaboration graphs, Attention-based mechanism with trainable parameters, and Personalized aggregation guided by collaboration graphs). We also performed experiments to analyze how FedAGHN uniquely captures collaborative relationships across different model layers, various stages of the FL process, and self-participation weight.
For future work, we suggest that whether there are more suitable input features is a problem worth exploring. In addition, the ability of capturing collaboration graphs is limited by the capacity of the hypernetwork, and the complexity of the hypernetwork and the learning cost would be a balance point worth careful study. Finally, FedAGHN is an aggregation optimization method that is decoupled from the optimization of the local update process, and combining it with other local update optimization methods is also a potential extension point.